### v1

# Improving Clinical Note Generation from Complex Doctor-Patient Conversation

Abstract
Writing clinical notes and documenting medical exams is a critical task for healthcare professionals, serving as a vital component of patient care documentation. However, manually writing these notes is time-consuming and can impact the amount of time clinicians can spend on direct patient interaction and other tasks. Consequently, the development of automated clinical note generation systems has emerged as a clinically meaningful area of research within AI for health. In this paper, we present three key contributions to the field of clinical note generation using large language models (LLMs). First, we introduce CliniKnote, a comprehensive dataset consisting of 1,200 complex doctor-patient conversations paired with their full clinical notes. This dataset, created and curated by medical experts with the help of modern neural networks, provides a valuable resource for training and evaluating models in clinical note generation tasks. Second, we propose the K-SOAP (Keyword, Subjective, Objective, Assessment, and Plan) note format, which enhances traditional SOAP (Podder, Lew, and Ghassemzadeh 2024) (Subjective, Objective, Assessment, and Plan) notes by adding a keyword section at the top, allowing for quick identification of essential information. Third, we develop an automatic pipeline to generate K-SOAP notes from doctor-patient conversations and benchmark various modern LLMs using various metrics. Our results demonstrate significant improvements in efficiency and performance compared to standard LLM finetuning methods.
Introduction
The generation of clinical notes is a critical task for healthcare professionals, serving as a vital component of patient care documentation. However, manually writing these notes is time-consuming, often taking clinicians between 10 to 30 minutes per note. This extensive time requirement may impact the amount of time clinicians can spend on direct patient interaction and other tasks. Consequently, the development of automated clinical note generation systems has emerged as an important area of research within AI for health. Recent advancements in large language models (LLMs) have significantly enhanced text summarization capabilities, offering the potential to streamline the creation of clinical notes. These models can assist clinicians in efficiently drafting their notes, allowing them to focus more on patient care and manage multiple tasks during consultations (Knoll et al. 2022).
Figure 1: Demonstration of the K-SOAP note generation pipeline
Despite this promise, the application in clinical note generation faces several challenges. Firstly, the quality of existing datasets is a major concern. While there are publicly available datasets like MTS-DIALOG  (Ben Abacha et al. 2023) and ACI-BENCH (wai Yim et al. 2023), they often fall short in terms of the complexity and length required to accurately simulate real-life clinical interactions. Secondly, the current format of clinical notes, such as SOAP notes, although clear and structured, can be time-consuming for clinicians to review, particularly as patient conditions become more complex. This increased complexity results in longer notes with dispersed critical information, complicating the decision-making process (Davidoff and Miglus 2011). Lastly, there is a notable lack of comprehensive benchmarking for LLM-based clinical note generation, hindering the evaluation and comparison of different models and approaches in this domain. Addressing these challenges requires finetuning LLMs with domain-specific knowledge, developing high-quality datasets that reflect real-life clinical dialogues, and establishing thorough benchmarking protocols to advance the efficacy and reliability of automated clinical note generation systems (Zhou et al. 2024).
Our work makes several significant contributions to the field of clinical note generation using large language models. First, we introduce CliniKnote, a comprehensive dataset of complex doctor-patient conversations paired with full clinical notes. Second, we propose the K-SOAP note format, enhancing traditional SOAP notes with a keyword section to improve information retrieval. Third, we perform thorough benchmarking of various large language models on this task and develop a novel method that significantly improves over standard LLM finetuning approaches.
All conversations in CliniKnote are simulated and recorded by medical experts to ensure they closely resemble real-life doctor-patient interactions, and the paired notes are also written by real medical experts. This dataset provides a valuable resource for training and evaluating models in clinical note generation tasks.
Using this dataset, we build the K-SOAP note format, which adds a keyword section at the top of traditional SOAP notes. This addition allows doctors to quickly recall domain knowledge at a glance. To achieve this, we combine clinical Named Entity Recognition (NER) and Relation Extraction (RE) to extract entities such as symptoms and diseases from the dialogue along with their relationships to the patient (e.g., present or absent). Previous clinical NER datasets like NCBI-disease (Dogan, Leaman, and Lu 2014) and BC5CDR (Li et al. 2016) only contain raw disease entities without relational information, highlighting the need for more comprehensive training resources like CliniKnote.
Furthermore, we finetune various modern large language models and develop an automatic pipeline to generate K-SOAP notes from the conversations. The workflow of this pipeline is shown in figure 1. We benchmark CliniKnote using various automatic evaluation to assess the quality of the generated notes and their practical applicability in real-life clinical settings. We also investigate the effects of data augmentation, finetuned domain knowledge, and different adapters for various section segmentations. We found that domain knowledge finetuned models with 15 different adapters for each section achieve the best performance. Our results demonstrate that our proposed method significantly reduces time complexity and enhances the performance of automatically generating K-SOAP notes. The time required for reviewing and correcting notes generated by this pipeline is significantly shorter than the original reviewing and writing time.
Related Work
Various methods have been developed for clinical note generation. For instance, Singh et al. (2023) build large-scale sequence-to-sequence models for generating clinical notes from patient-doctor conversations. The use of pre-trained large language models (LLMs) for this task is increasingly popular (Zhou et al. 2024). Biswas and Talukdar (2024) demonstrates advanced prompting techniques to generate draft clinical notes using LLMs, while Grambow, Zhang, and Schaaf (2022) introduces in-domain pre-training to enhance transformer models’ clinical summarization performance. Additionally, the MEDIQA-2023 Dialogue2Note shared tasks (Tang et al. 2023) include finetuning pre-trained dialogue summarization models and using few-shot in-context learning with GPT-4 to address medical dialogue summarization challenges.
For publicly available clinical conversation-note datasets, there are several notable examples. Dr.Summarize (Joshi et al. 2020), generated from a telemedicine platform, and the dataset presented by Chintagunta et al. (2021), created using GPT-3, both follow a snippet-summary format. For well-structured SOAP medical notes, ACI-BENCH (wai Yim et al. 2023) and PriMock57 (Korfiatis et al. 2022) feature role-played dialogues with complete SOAP notes but contain only 207 and 57 data points, respectively, which is insufficient for training modern large language models. MTS-DIALOG (Ben Abacha et al. 2023), previously the largest dataset with clinical notes, includes 1,700 doctor-patient conversations (16k turns and 18k sentences) and summarized clinical notes (6k sentences). However, MTS-DIALOG consists of segmented snippets and paired notes that include only some sections rather than full notes.
Named Entity Recognition (NER) and Relation Extraction (RE) are essential for extracting structured information from unstructured text in natural language processing. Recent advancements in deep learning, particularly with BiLSTM-CRF and transformer-based models like BERT, have significantly improved accuracy in identifying entities across various domains, including the medical field. Notably, recent work combining NER and RE for extracting complex information from scientific texts (Dunn et al. 2022) and utilizing label-supervised Llama finetuning (Li et al. 2023) has achieved state-of-the-art performance. Building on these advancements, our work integrates cutting-edge NER and RE techniques with large language models to enhance clinical note generation from doctor-patient dialogues.
As large language models scale up, finetuning and storing all the parameters is prohibitively costly and eventually becomes practically infeasible. Recent advancements in natural language processing have led to the development of parameter-efficient pre-trained large language models. These models aim to reduce the computational and memory requirements while maintaining or even improving performance across various tasks. One notable approach is the use of adapter modules, which allow for the finetuning of a small number of additional parameters while keeping the majority of the pre-trained model fixed (Houlsby et al. 2019). This technique has been proven effective in domain adaptation and multilingual settings. Additionally, methods such as LoRA (Low-Rank Adaptation) (Hu et al. 2021) and Qlora(Quantized Low-Rank Adaptation) (Dettmers et al. 2023) have shown promise in optimizing model efficiency by introducing minimal changes to the original model architecture. These innovations are particularly relevant for applications in the medical domain, where models need to handle large volumes of complex data without incurring prohibitive computational costs. Our work builds upon these foundations by integrating parameter-efficient techniques with state-of-the-art summarization and information extraction models to enhance the generation of clinical notes from doctor-patient dialogues.
Figure 2: Keyword extraction in clinical dialogue.
CliniKnote Dataset Assessment
In this section, we discusse how we synthesize the conversations, SOAP notes, and keyword sections to create our CliniKnote dataset. We also highlight CliniKnote’s advantages in terms of scale and structure by comparing with the previous largest public available note dataset MTS-DIALOGUE (Ben Abacha et al. 2023). The appendix A shows a full example of CliniKnote.
Data Synthesis
CliniKnote’s dialogue and SOAP note sections are curated in a similar process as described in  (Fareez et al. 2022). A team of composed of senior medical students, medical residents, and medical doctors conducted simulated medical conversations. Following the completion of each simulated medical exam, the individual assuming the role of the physician would write a standard SOAP clinical note summarizing the findings, assessment, and medical plan of the simulated visit. Because each exam was simulated, there is no personal health information included in the dataset. The audio of the conversation was initially transcribed using automatic speech recognition (ASR) and then manually validated and corrected to minimize word error rates (WER). Through manual validation the WER of the transcripts were estimated to be approximately 1%.
CliniKnote’s keyword section contains all the symptoms and diseases mentioned in the conversation together with their relations to the patients labeled by GPT-4. Each keyword is prefixed to indicate its relation to the patient. Through the keyword section, we can directly find the useful symptom and disease information without looking at the full medical note. We have 8 types of entities in keyword sections in total, which are PRESENT SYMPTOM, PRESENT DISEASE, ABSENT SYMPTOM, ABSENT DISEASE, FAMILY DISEASE, PAST DISEASE, UNKNOWN DISEASE and UNKNOWN SYMPTOM respectively. The prefix PRESENT and ABSENT describe the relation between this entity and the patient. The prefix FAMILY and PAST mean that the patient has a family history of this entity while the prefix UNKNOWN means we are not sure this entity is present or absent currently and some further tests may need to be done. Figure 2 shows a demo of keyword section.
CliniKnote MTS-DIALOG
Dialogue Summary Dialogue Summary
Turns Sentences Words Sentences Words Turns Sentences Words Sentences Words
count 70,557 117,131 1,144,432 54,571 586,052 15,969 18,406 241,685 5,870 81,299
mean 59 98 954 45 490 9 11 142 3 48
max 102 155 1,403 70 796 103 136 1,951 57 1,182
min 40 50 700 10 371 - - - - -
Table 1: Data statistics of CliniKnote and MTS-DIALOG.
Data Augmenting
During the development of medical exams, the simulations covered a limited number of allergies, thus the dataset lack diversity of the allergy field. Using vanilla CliniKnote for note generation might cause identical allergy generation ignoring the specific patient input. In order to enhance the data diversity of CliniKnote, we augment the allergy section of CliniKnote by randomly sampling more allergies. To be more specific, we first generate a list of 30 common allergies and pick a number from 0 to 4 randomly. Then we randomly pick that number of allergies from the list, feed those allergies and the original dialogue into GPT-4, and prompt GPT-4 to replace the original allergies with the allergies we feed. The prompt is shown in Appendix A After replacing the allergies, we do a string comparison to make sure that is the only modified part in the dialogue. Through data augmentation for the allergy part, we have a more complementary version of CliniKnote, which could be generalized for more clinical-related tasks.
Entity Type Count
PRESENT SYMPTOM 12034
PRESENT DISEASE 2442
ABSENT SYMPTOM 24687
ABSENT DISEASE 283
FAMILY DISEASE 1922
PAST DISEASE 1001
UNKNOWN DISEASE 1538
UNKNOWN SYMPTOM 1270
Table 2: Entity counts across different categories.
Data Statistics
The final CliniKnote we introduced has 1000 dialogues in training set and 200 dialogues in test set, both paired with the proposed K-SOAP medical notes including the sections of Chief Complaint, History Of Presenting Illness, Past Medical History, Past Surgical History, Family History, Allergies, Social History, Medication, Immunization History, Review of Systems, Vital Signs, Physical Exam, Diagnostic, Assessment, Plan, and keywords. In training set, We have 952,316 words and 489,594 words in medical notes. In testing set, we have 192,116 words in dialogues and 96,458 words in medical notes.
Compared with MTS-DIALOG, we have more numerous data points in CliniKnote and more complex dialogues that are closer to real-life clinical conversations. Table 1 shows the comparison between our CliniKnote and MTS-DIALOG, which is the largest publicly available dataset of medical conversations and associated notes, demonstrating our large scales in word count and turns. Our word count is 5 times greater than that of MTS-DIALOG in dialogues and 7 times larger in summary. The final keyword section contains 12,286 short dialogues segmented from the original dialogues paired with the entities with relation labeled. Each short dialogue has 6 turns in total so that it will be easier for NER models to process. Table 2 shows the specific count for each type of entity.
K-SOAP Generation Based On Medical Dialogue
The complete task aims to generate K-SOAP note based on the input dialogue. We divide this task into two parts. The first part is to generate the traditional SOAP note and the second is to do keyword extraction in NER format. The concatenation of these two parts made up to the final K-SOAP note.
SOAP Clinical Note Generation
The main task for our CliniKnote dataset is to generate the traditional SOAP clinical notes. To compare the performance of models with different sizes, domain knowledge, and released versions, we selected the following publicly available models to finetune for our benchmarking: qCammel-13b (Toma 2023), a finetuned version of the Llama-2 13B model, trained on a distilled dataset of 15,000 medical domain instructions using QLoRA; Llama2-13b (Touvron et al. 2023); Llama3-8b (AI@Meta 2024); and OpenBioLLM-Llama3-8B (Ankit Pal 2024), which is finetuned on the biomedical domain. Additionally, we include commercial models such as GPT-3.5-turbo and GPT-4 for our benchmarking. We use offical QLoRA (Dettmers et al. 2023) to finetune all the large language models. For this supervised learning part, the input is the prompt filled with the original dialogue and the output is the clinical note in CliniKnote dataset reformatted in json mode as it will be more convenient to access. After finetuning, the Parameter-Efficient finetuning (PEFT) adapters (Mangrulkar et al. 2022) will be saved and loaded to do further evaluation.
Model Rouge-1 Rouge-2 Rouge-L
Rouge-
lsum
Bertscore-P Bertscore-R Bertscore-F1 Bleurt QuestEval
baseline models
Bart-full
0.44 0.29 0.33 0.417 0.908 0.808 0.853 0.563 0.498
Bart-section-4
0.633 0.388 0.459 0.606 0.869 0.887 0.878 0.518 0.463
Bart-section-15
0.591 0.367 0.406 0.572 0.801 0.842 0.821 0.372 0.484
biobart-full
0.439 0.303 0.336 0.417 0.906 0.810 0.855 0.567 0.506
biobart-section-4
0.633 0.402 0.382 0.608 0.867 0.890 0.878 0.512 0.459
biobart-section-15
0.583 0.366 0.399 0.566 0.784 0.837 0.809 0.318 0.478
open-source LLMs
Llama2-13b-full
0.641 0.409 0.505 0.591 0.871 0.852 0.861 0.534 0.461
Llama2-13b-section-4
0.678 0.437 0.516 0.632 0.871 0.852 0.861 0.553 0.399
Llama2-13b-section-15
0.656 0.426 0.506 0.613 0.872 0.861 0.867 0.495 0.426
qCammel-13b-full
0.639 0.395 0.477 0.568 0.879 0.865 0.872 0.581 0.475
qCammel-13b-section-4
0.465 0.289 0.354 0.426 0.889 0.866 0.877 0.476 0.327
qCammel-13b-section-15
0.692 0.467 0.538 0.646 0.896 0.893 0.895 0.474 0.395
Llama3-8b-full
0.641 0.409 0.505 0.591 0.893 0.878 0.885 0.595 0.476
Llama3-8b-section-4
0.678 0.437 0.516 0.632 0.867 0.860 0.864 0.548 0.468
Llama3-8b-section-15
0.656 0.426 0.506 0.613 0.866 0.857 0.862 0.550 0.465
OpenBioLLM-Llama3
-8B-full
0.372 0.189 0.254 0.307 0.839 0.798 0.818 0.421 0.421
OpenBioLLM-Llama3
-8B-section-4
0.386 0.194 0.253 0.340 0.834 0.816 0.825 0.319 0.364
OpenBioLLM-Llama3-8B
-section-15
0.298 0.207 0.248 0.282 0.834 0.816 0.825 0.297 0.316
Commercial models
GPT-3.5-Turbo-
zeroshot
0.592 0.322 0.441 0.562 0.802 0.846 0.823 0.500 0.407
GPT-3.5-Turbo-
oneshot
0.705 0.451 0.580 0.663 0.855 0.889 0.871 0.560 0.441
GPT-4o-
zeroshot
0.630 0.363 0.491 0.602 0.816 0.847 0.831 0.467 0.405
GPT-4o-
oneshot
0.753 0.519 0.634 0.704 0.878 0.899 0.889 0.610 0.493
Table 3: Overall results for medical dialogue to note task on CliniKnote dataset.
Keyword Extraction
Besides generating SOAP content, we further generate the keyword section for each clinical dialogue to reduce the reviewing complexity. We treat this task as a NER task but with relation as the prefix. For keyword extraction, we use Label-supervised unmasked Llama finetuning (Li et al. 2023). Instead of autoregressive generation, we add a label predictor at the top of a Llama-7b model. The label predictor is a one-layer linear feed forward neural network. The dialogue sequences decoded by Llama will be fed into this predictor to predict the labels. We also keep the idea of removing the causal masks because their empirical studies show that using token representations learned with causal masks significantly underperforms in token classification tasks. In order to make this task more understandable to the model, we merge the entities into only 4 types, which are PRESENT, ABSENT, FAMILY and UNKNOWN. To be specific, we set the word to tag dictionary as ‘O’: 0, ‘B-PRESENT’: 1, ‘B-ABSENT’: 2, ‘B-FAMILY’: 3, ‘B-UNKNOWN’: 4, ‘I-PRESENT’: 5, ‘I-ABSENT’:6, ‘I-FAMILY’:7, ‘I-UNKNOWN’:8 while B means start and I means inside. The model will determine whether a word is a symptom or disease and what is the relation with the patient at the same time.
Evaluation of K-SOAP Note Generation
We present various evaluation metrics on the benchmarking of CliniKnote, to investigate the generated notes’ quality and time complexity reduction. We aim to demonstrate the effectiveness of CliniKnote in producing high-quality clinical notes and its potential to enhance the workflow in medical practice by saving time and improving accuracy.
Evaluation Metrics
The evaluation of clinical note generation is particularly challenging, as it is difficult to accurately measure the relevance and quality of the generated notes due to the complexity and specificity of medical information.
For keyword evaluation, we only record and compare the accuracy, precision, recall and F1 score. For SOAP part note evaluation, we employ several metrics to ensure a comprehensive assessment. We start with the most traditional metric, ROUGE (Lin 2004), which evaluates similarity based on common subsequences between the generated and reference texts. In addition to ROUGE, we introduce more advanced metrics: BERTScore (Zhang* et al. 2020) and BLEURT  (Sellam, Das, and Parikh 2020)  (Pu et al. 2021). These metrics leverage contextual embeddings from BERT to evaluate the semantic similarity between generated and reference texts. BERTScore uses token embeddings to compute precision, recall, and F1 scores, providing a more nuanced and context-aware evaluation compared to traditional metrics. Similarly, BLEURT incorporates pre-trained transformer models to better capture the meaning and fluency of the text, offering insights into the semantic quality of the generated notes. We use Roberta-large model for BERTScore metric and the newest checkpoint BLEURT-20 for BLEURT metric.
Finally, we include QuestEval (Scialom et al. 2021) in our evaluation framework. Unlike traditional metrics, QuestEval does not require any ground truth references. Compared with ROUGE, QuestEval substantially improves the correlation with human judgments over four evaluation dimensions: consistency, coherence, fluency, and relevance. By incorporating question-answering models, QuestEval evaluates the generated text’s ability to provide accurate and relevant information, making it particularly suitable for complex domains like clinical note generation.
Model Acc F1 Pre Recall Eval_Loss
biobert 0.933 0.735 0.755 0.715 0.197
scibert 0.921 0.669 0.723 0.622 0.221
pubmedbert 0.919 0.685 0.689 0.681 0.230
bluebert 0.903 0.554 0.679 0.467 0.285
bert 0.915 0.653 0.677 0.630 0.242
Llama2-7b-in-context - 0.692 0.391 0.500 -
Llama2-7b-QA - 0.642 0.731 0.657 0.307
Llama2-7b-
label-supervised 0.971 0.752 0.726 0.780 0.107
Table 4: Keyword performance comparison of different models.
Generated Section Reference Section
1 ”Chief Complaint”: ”Skin rash on arms and legs.” ”Chief Complaint”: ”Skin issues on forearm.”
2 ”History Of Presenting Illness”: ”62-year-old female presents with a skin issue that started as a small red patch on her forearm about a month ago. The patch has been getting larger and itchier, rated as a 7 out of 10 in severity. Cold compresses provide some relief, while heat and wearing long sleeves aggravate the condition. The itching is affecting her sleep and self-esteem.” ”History Of Presenting Illness”: ”62-year-old female presents with a skin issue that started about a month ago. Initially a small red patch on the forearm, which has been getting larger and itchier. Itchiness rates at about a 7 out of 10. Symptoms worsen at night and when wearing long sleeves. Cold compresses provide slight relief, while heat and fabric contact aggravate the condition. The issue has impacted sleep and caused self-consciousness about appearance. No prior occurrence of similar symptoms. Over-the-counter creams have been tried with little effect. ”
3 ’Family History’: [’Father had psoriasis, mother has a history of eczema. No other major illnesses that I’m aware of.’] ’Family History’: [’Father had psoriasis, mother has eczema.’]
4 ’Allergies’: [’Latex: severe skin irritation’] ’Allergies’: [’Latex - causes severe skin irritation.’]
5 ”Social History”: ”Patient is a retired librarian, widowed with two adult children. She enjoys gardening and has never smoked. Drinks wine occasionally with dinner.” ”Social History”: ”Retired librarian, widowed with two adult children. Enjoys gardening, non-smoker, minimal alcohol consumption.”
Table 5: Examples of generated summaries by qCammel-13b-section-15.
Experiments
In this section, we focus on the performance of various models after finetuning on CliniKnote. We also discuss different finetuning methods for keyword extraction.
Experimental Settings
We train all the models mentioned in this benchmarking with one A100-80GB GPU. For clinical note generation task, we train all the large language models for two epochs. We use the default setting for question-answer finetuning with QLoRA. We set the learning rate to 0.0002 and the warmup ratio to 0.03. For QLoRA hyperparameters, we set LoRA $r$ to 64 and the loRA $\alpha$ to 16 and all the models are quantized in 4 bits. For keyword task, we just use LoRA without quantizing the model. We set batch size = 8, learning rate = 1e-4, Max length = 256, LoRA $r$ = 256, LoRA $\alpha$=768, LoRA dropout=0.1. We train Llama2-7b model for 5 epochs.
Clinical Note Generation Results
In this section, all models are finetuned and tested in 3 modes, which are full note mode, section-4 mode and section-15 mode that contain different numbers of LoRA adapter (Hu et al. 2021). Full note mode means we only train one adapter to generation the full clinical note. Section-4 mode means we train four different adapters to generate the full note. How we divide the four parts is a little different from how we separate Subjective, Objective, Assessment, and Plan parts. We only have Chief Complaint and History Of Presenting Illness in the first part. The second part contains Past Medical History, Past Surgical History, Family History, Allergies, Social History, Medication List, Immunization History and Review of Systems. The third part contains Vital Signs and Physical Exam while the final part contains Assessment, Diagnosis and Plans. Section-15 mode means we train 15 different adapters to generate all the fifteen sections contained in the note separately. As for commercial models, we try full note zeroshot generation and oneshot(one demo in prompt) in-context learning.
For baseline models, we choose Bart and Biobert as baseline models. Bidirectional autoregressive transformer Bart is the SOTA transformer model for note generation. For matching clinical tasks, we use the continued pre-training on PubMed abstract model of BART-large version. We further compare the performance of BioBart, which has the same model structure as BART but possesses distinct tokenizers and vocab- ulary size.
Dialogue D: When did you first notice this rash? P: It started about two weeks ago, just a small patch. D: Has anything in particular seemed to make it better or worse? P: It gets itchier at night, and when I sweat.
Keywords [‘1rash’, ‘—’, ‘1itchier’, ‘5at’, ‘5night’, ‘—’, ‘5sweat’]
[‘1rash’, ‘—’, ‘1itchier’, ‘5at’, ‘5night’]
Dialogue D: Any cough or shortness of breath? P: No. D: Any changes in urination or blood in the urine? P: No. D: Any changes in bowel movements, nausea, vomiting, or abdominal pain? P: No.
Keywords [‘2cough’, ‘—’, ‘2shortness’, ‘6of’, ‘6breath’, ‘—’, ‘2changes’, ‘6in’, ‘6urination’, ‘—’, ‘2blood’, ‘6in’, ‘6the’, ‘6urine’, ‘—’, ‘2changes’, ‘6in’, ‘6bowel’, ‘6movements’, ‘—’, ‘2nausea’, ‘—’, ‘2vomiting’, ‘—’, ‘2abdominal’, ‘6pain’]
[‘2cough’, ‘—’, ‘2shortness’, ‘6of’, ‘6breath’, ‘—’, ‘2changes’, ‘6in’, ‘6urination’, ‘—’, ‘2blood’, ‘6in’, ‘6the’, ‘6urine’, ‘—’, ‘2changes’, ‘6in’, ‘6bowel’, ‘6movements’, ‘—’, ‘2nausea’, ‘—’, ‘2vomiting’, ‘—’, ‘2abdominal’, ‘6pain’]
Table 6: Comparison of Keywords Extraction in Dialogues: Red for Prediction, Blue for Ground Truth.
Table 3 shows the performance of all the models with the automatic metrics we pick. qCammel-13b-section-15 achieves overall best performance on ROUGE and BERTScores among all open-scouce models while GPT-4o-oneshot is considered the best among all commercial models. To our surprise, section-15 mode does not show significant enhancement on the performance even though it takes 11 and 14 more adapters than other modes. All three modes of Llama3-8b show decent results. The section-4 mode of Llama3-8b shows second best peformance on ROUGE score and full mode has highest score on Bleurt, QuestEval and second highest score on Bertscore among all open-source models. However, OpenBioLLM-Llama3-8b has poor performance on almost all metrics. All three modes of OpenBioLLM-Llama3-8b shows lowest scores on ROUGE, BertScore, Bleurt and QuestEval, even lower than baseline models. The domain knowledge insertion triggers decreasing on information extraction ability. The benchmarking shows that appropriate in-domain instruction is beneficial while it could also jeopardize the model’s original extraction ability. Table 5 shows some sections generated by qCammel-13b-section-15 and compares them with the reference sections. After data augmenting, the occurrence rate of penicillin as an allergy drops by 11.5% The actual occurrence rate of penicillin is 19.5% in test set. However, the qCammel-13b-section-15 model predicts penicillin as an allergy 31.5% of the time without data augmentation. Specifically, the model tends to output penicillin when no allergy is mentioned in the dialogue. After applying data augmentation, the penicillin prediction rate drops to 20%, which aligns much closer to the real distribution.
Keyword Extraction Results
We use BioBart, SciBert, PubMed, BlueBert and Bert as baseline models for key words extraction task. For baseline models, we finetune the pretrained models on CliniKnote dataset. The learning rate is set to 1e-4, batch size is set to 128 and we train all baseline models for 10 epochs. Besides the label-supervised Llama finetuning, we also tried in-context learning and question-answer based finetuning on the same Llama2-7b model. Table 4 shows the results among all the models and Llama2-7b with our label-supervised finetuning dominates all the metrics except for Precision and has the lowest $EvalLoss$. The precision of Llama2-7b-label-supervised is slightly lower than that of biobert. Neither in-context learning nor QA-based finetuning on Llama2-7b has decent performance on this task since it combines RE as prefix and makes it harder to predict for autoregressive LLM. Table 6 shows two demos of the Llama2-7b-label-supervised’s prediction and the ground truth. Label-supervised finetuning enhances the entity recognition ability of large language models while maintaining their comprehension skills. The results also demonstrate that raw llms cannot handle this keyword extraction task with decent performance, which indicates the necessity of the training resources provided in CliniKnote keyword section.
Conclusion
In this study, we addressed the critical task of automatic clinical note generation to enhance the efficiency and accuracy of medical documentation. We introduced the K-SOAP format, which includes keywords at the top of traditional SOAP notes for quicker review and comprehension. Additionally, we presented CliniKnote, a large-scale dataset containing complex dialogues paired with K-SOAP notes. Training large language models on CliniKnote demonstrated substantial improvements in generating accurate clinical notes. Our comprehensive evaluation ensured a thorough assessment of the generated notes, making the assessment of the quality of generated notes more straightforward and robust. This work advances clinical note generation, driving further research and saving doctors time in documentation tasks.
Limitations
Although CliniKnote is created by experts to simulate real-life clinical inquiries, there are still robustness issues as 100% of the data in our dataset are clean. In real-life conversations, ASR systems may not always produce clean transcripts due to various factors such as accents, interjections, and expression variations. These factors pose significant challenges for the robustness of clinical note generation.
Additionally, the range of models used in this benchmarking is relatively narrow. Many open-source models with larger sizes or better configurations may be suitable for generating K-SOAP notes, indicating significant potential for improvement in this pipeline.
Finally, the current evaluation metrics do not perfectly reflect the quality of clinical notes, as they struggle to capture key information and crucial errors in the medical domain. We hope that more advanced evaluation metrics for text summarization and note generation will be developed to address this issue. Moreover, human-expert evaluation still need to be completed. We are planning a more extensive evaluation for both CliniKnote and real-life clinical conversations with finetuned models to assess accuracy and the time savings for doctors in real clinical scenarios.
Ethics Statement
No protected health information were used in the creation of this dataset.
Appendix A Example Appendix
K-SOAP Demo
Dialogue D: Good day! I’m Dr. Keir MacDonald, and I’ll be looking after your health concerns today. I see here you’re Mr. Callahan MacLean, correct? How are you feeling this morning?
P: Morning, Dr. MacDonald. I’ve been better, to be honest. I’ve been having some trouble with frequent urination and a burning sensation when I do. It’s been going on for a few days now.
D: I’m sorry to hear that, Mr. MacLean. Can you tell me more about the onset of these symptoms? Have they been constant or intermittent? And on a scale from 1 to 10, how would you rate any pain associated with urination?
P: It started suddenly about four days ago. The discomfort is pretty constant, especially the urge to urinate. The pain is around a 4, but it’s more the frequency that’s troubling me. It’s affecting my work since I have to run to the bathroom so often.
D: Understood. Have you experienced these symptoms before, or is this the first time?
P: This is the first time, and I haven’t had any tests done for it yet.
D: Let’s delve into your medical history a bit. Aside from this, any other chronic conditions or past illnesses I should be aware of?
P: I’ve been managing type 2 diabetes for the past decade, and I have a history of kidney stones.
D: Any allergies, Mr. MacLean?
P: Just penicillin.
D: Are you up to date with your vaccinations, including your flu shot?
P: I am. Got my flu shot last month.
D: How about family medical history? Any similar issues or other conditions in your family?
P: My father had issues with his prostate, and my mother has a history of urinary tract infections.
D: And what about your personal life? Do you live with anyone, are you currently working, and what do you do for leisure?
P: I’m divorced, live alone, and I’m a retired librarian. I enjoy reading and bird watching. I don’t smoke and I’ll have the occasional beer, maybe once a month.
D: Let’s talk about your current medications. What are you taking, and what dosages?
P: I’m on metformin 500 mg twice daily for my diabetes, and that’s about it.
D: Now, I’ll go through a review of systems. Any fever, chills, or recent weight loss?
P: No fever or chills, and my weight has been stable.
D: Any congestion, sore throat, or ear pain?
P: No, nothing like that.
D: Chest pain or palpitations?
P: No, my chest’s been fine.
D: Cough or difficulty breathing?
P: No respiratory issues.
D: Any changes in bowel movements or abdominal pain?
P: No, my bowels have been normal.
D: Dizziness, headaches, or memory issues?
P: No neurological symptoms.
D: Joint pains or stiffness?
P: My knees ache occasionally, but nothing significant.
D: Any intolerance to heat or cold?
P: No, I’m comfortable with the temperature.
D: I see. Let’s proceed with the physical exam. You don’t appear to be in any respiratory distress. Let’s check your heart rate… It’s 78 beats per minute. Blood pressure is 130/85, and respiratory rate is 14 breaths per minute. All within normal ranges.
D: Your affect is calm and you’re well-oriented. No redness or swelling in the head and neck region. Pupils are equal and reactive to light. Heart sounds reveal normal S1 and S2 without murmurs. Lungs have good air entry bilaterally, no wheezing or other sounds. Abdomen is soft, non-tender, with normoactive bowel sounds. Cognitive screening is normal.
P: That’s good to hear, doctor.
D: Based on your symptoms and history, it seems like you may have a urinary tract infection. We’ll need to run a urinalysis to confirm and possibly a culture to identify the bacteria. I’ll prescribe a course of antibiotics that’s safe for someone with a penicillin allergy. We should also monitor your blood sugar closely, as infections can affect glucose levels.
P: Sounds like a plan, Dr. MacDonald.
D: I’ll give you a requisition for the lab work. Try to get it done as soon as possible. In the meantime, increase your water intake to help flush out your system. I’d like to see you back here in a week to review the lab results and check on your progress. If your symptoms worsen or you develop a fever, please go to the emergency room immediately. Do you have any other concerns or questions for me?
P: No, that covers everything. Thank you for your help.
D: You’re welcome, Mr. MacLean. Take care and I’ll see you soon.
Note: Figure LABEL:fig:full_demo shows the corresponding KSOAP note.
Allgeries Augmenting Prompt
allergens = [ ”Milk”, ”Eggs”, ”Fish”, ”Shellfish”, ”Tree nuts”, ”Peanuts”, ”Wheat”, ”Soybeans”, ”Sesame seeds”, ”Mustard”, ”Sulfites”, ”Lupin”, ”Almonds”, ”Cashews”, ”Walnuts”, ”Pistachios”, ”Hazelnuts”, ”Brazil nuts”, ”Pecans”, ”Macadamia nuts”, ”Pine nuts”, ”Gluten”, ”Oats”, ”Rye”, ”Barley”, ”Corn”, ”Kiwi”, ”Banana”, ”Avocado”, ”Strawberry”, ”Citrus fruits”, ”Apples”, ”Cherries”, ”Peaches”, ”Tomatoes”, ”Carrots”, ”Celery”, ”Garlic”, ”Onions”, ”Chili peppers”, ”Peppercorns”, ”Cinnamon”, ”Cloves”, ”Vanilla”, ”Ginger”, ”Anise seeds”, ”Fenugreek”, ”Sunflower seeds”, ”Poppy seeds”, ”Mollusks”, ”Crustaceans”, ”Artificial colors”, ”Artificial flavors”, ”Monosodium glutamate (MSG)”, ”Aspartame”, ”Nutrasweet”, ”Saccharin”, ”Soy lecithin”, ”Soy protein”, ”Whey protein”, ”Casein”, ”Yeast”, ”Tapioca”, ”Seaweed”, ”Agar agar”, ”Mango”, ”Penicillin”, ”Amoxicillin”, ”Cephalexin”, ”Clindamycin”, ”Vancomycin”, ”Sulfa drugs (Sulfonamides)”, ”Tetracycline”, ”Doxycycline”, ”Aspirin” ]
Prompt = You are a data augmenting assistant and you job is to replace the allergies in the following dialogue between a patient and a doctor. The doctor may ask the patient’s allergies and you should change the patient’s answer with one or more random allergens. If the doctor didn’t ask about the allergies, just output the original dialogue and don’t modify anything. The allergies you need to replace with is ”’+allergens+”’ Please organize the output in json format:{{’output’: new dialogue}} or {’output’: ’not modified’} if doctor doesn’t ask in the dialogue.
Dialogue:”’+dialogue